# Cutoff for Contingency Table Random Walks

### v1

@@ 2.2 Technical Lemmas
We break the sum into two parts and bound them respectively as follows. We can write
@@
For any $n\in\mathbb{N}$, we can write
@@
$\frac{1}{n^{2}}\sum_{\ell=1}^{\lfloor n^{1/4}\rfloor}\sum_{k=1}^{\lfloor n^{1/4}\rfloor}n^{\frac{2(k+\ell+1)}{(k+1)(\ell+1)}}\leq\frac{1}{n^{2}}\cdot n^{1/4}\cdot n^{1/4}\cdot n^{3/2}=1.$
@@
$\frac{1}{n^{2}}\sum_{\ell=\lfloor n^{1/4}\rfloor}^{n}\sum_{k=\lfloor n^{1/4}\rfloor}^{n}n^{\frac{2(k+\ell+1)}{(k+1)(\ell+1)}}\leq\frac{1}{n^{2}}\sum_{\ell=\lfloor n^{1/4}\rfloor}^{n}\sum_{k=\lfloor n^{1/4}\rfloor}^{n}n^{\frac{4(k+\ell)}{\ell k}}\leq\frac{1}{n^{2}}\cdot n\cdot n\cdot n^{8n^{-1/4}}\leq e^{32/e}$
as $x^{8x^{-1/4}}$ is maximized as $e^{32/e}$ at $x=e^{4}$ by elementary calculus. Now it suffices to bound the sum for $\ell\leq n^{1/4}$ and $k\geq n^{1/4}$, and the case $\ell\geq n^{1/4}$ and $k\leq n^{1/4}$ would follow in an analogous way. We consider the scenario of $\ell=1$ separately. When $\ell=1$, we have
$\frac{1}{n^{2}}\sum_{k=\lfloor n^{1/4}\rfloor}^{n}n^{\frac{k+2}{k+1}}=\frac{1}{n}\sum_{k=\lfloor n^{1/4}\rfloor}^{n}n^{\frac{1}{k+1}}\leq\frac{1}{n}\cdot n\cdot n^{n^{-1/4}}\leq e^{4/e}.$
@@
$\frac{1}{n^{2}}\sum_{\ell=2}^{\lfloor n^{1/4}\rfloor}\sum_{k=\lfloor n^{1/4}\rfloor}^{n}n^{\frac{2(k+\ell+1)}{(k+1)(\ell+1)}}\leq\frac{1}{n^{2}}\cdot n^{1/4}\cdot n\cdot n^{\frac{2(n^{1/4}+3)}{3n^{1/4}}}=n^{-1/12}\cdot n^{2n^{-1/4}}\leq e^{8/e}.$
Therefore, we conclude that
@@
Let $F(x):=\sum_{j=-N}^{N}\exp(-c(j-x)^{2})$, where $N\in{\mathbb{N}}$ and $c>0$ are constants. Then for all $c$ large enough (independent of $N$), $F(x)$ attains its global maximum at $x=0$.
@@
First we observe that for any $x\in(-1/2,1/2]$, we have
@@
by comparing the summands in $F(x+k)$ and $F(x+k+1)$, $k\in{\mathbb{Z}}$. Thus it suffices to maximize $F$ on the interval $(-1/2,1/2]$.
@@
We will only consider large $c>0$. If $x\not\in\bigcup_{j=-N}^{N}B_{j}(1/\sqrt{2c})$, then $F^{\prime\prime}(x)$ is necessarily positive since all the summands above are positive. Combining this with the previous observation, we only need to maximize $F$ on the interval $[-1/\sqrt{2c},1/\sqrt{2c}]$.
@@
Note that for all $c$ large and any fixed $x\in[0,1/\sqrt{2c}]$, the function
@@
For all large $c$, the function $g(x):=(1+x+\frac{1}{2c})e^{-c(1+x)^{2}}$ is convex on the interval $[-1/\sqrt{2c},1/\sqrt{2c}]$. So for $x\in[0,1/\sqrt{2c}]$,
@@
Now it remains to show that
@@ 3.1 Setup
We first provide a general method for mixing time lower bounds for a family of random walks on the torus ${\mathbb{Z}}_{q}^{n}$. We parameterize the torus ${\mathbb{Z}}_{q}^{n}$ by $({\mathbb{Z}}\cap[-q/2,q/2))^{n}$, i.e. these are the representatives we will use for elements in ${\mathbb{Z}}_{q}^{n}$.
Let $\{Y_{m}\}_{m\in{\mathbb{N}}}$ be an arbitrary irreducible symmetric discrete-time Markov chain on ${\mathbb{Z}}^{n}_{q}$ with transition matrix $P$, such that ${\mathbb{P}}(Y_{m}=y_{m}|Y_{m-1}=y_{m-1})=\mu(y_{m}-y_{m-1})$. Suppose that $\|Y\|_{1}\leq r$ for some real number $r$ in the $L_{1}$ norm, and $\|Y\|_{\infty}<q/2$ in the sup norm. We set $Y_{0}=0$. Symmetry of $\{Y_{m}\}$ implies $\mu(y)=\mu(-y)$ for all $y\in{\mathbb{Z}}_{q}^{n}$. Let $Y\in{\mathbb{Z}}^{n}$ be a lattice-valued random vector with law $\mu$ mapped to $({\mathbb{Z}}\cap[-q/2,q/2))^{n}$. We will assume $Y$ has the same marginal variance $\dfrac{\sigma^{2}}{n}$ for each coordinate. Write $\operatorname{Var}(Y)=\dfrac{\sigma^{2}}{n}\Gamma$ where $\Gamma$ is the correlation matrix of $Y$. Let $\{X_{t}\}_{t\geq 0}$ be the rate-1 continuous-time Markov chain on ${\mathbb{Z}}_{q}^{n}$ with rate matrix $Q=P-I$ and transition kernel $H_{t}=\exp(tQ)$. Set $X_{0}=0$.
@@ 3.2 Representations and Eigenvalues
By Theorem 3 of [10] (a more detailed proof could be found in Theorem 6 of [6]), the Fourier transforms $\{\hat{\mu}(\rho_{y}):y\in{\mathbb{Z}}_{q}^{n}\}$ are precisely the eigenvalues of the discrete-time process $\{Y_{m}\}$, and the characters are the eigenfunctions of $P$. Thus the eigenvalues $\{\lambda_{j}:1\leq j\leq q^{n}\}$ of the continuous-time process $\{X_{t}\}$ are
$\{\lambda_{j}:1\leq j\leq q^{n}\}=\{\exp(t(\Phi_{Y}(2\pi y/q)-1)):y\in({\mathbb{Z}}\cap[-q/2,q/2))^{n}\}.$
@@ 3.3.1 Forming a Distinguishing Statistic
Let $\psi_{i}(x):=\cos\left(\dfrac{2\pi x_{i}}{q}\right)$ for $({\mathbb{Z}}\cap[-q/2,q/2))^{n}$. We form a distinguishing statistic
@@
For notation, let $\delta_{i}$ be the vector whose $i$-th coordinate is $1$ and zero elsewhere. We know that each $\psi_{i}$ is an eigenfunction of $H_{t}$ with a corresponding eigenvalue of
@@
By the fourth-order Taylor expansion of the characteristic function $\Phi_{Y}$, we obtain that for $i\neq j$,
$\exp\left(-\frac{2\pi^{2}t\sigma^{2}}{nq^{2}}\right)\leq\exp\left(t\left(\Phi_{Y}\left(\frac{2\pi\delta_{i}}{q}\right)-1\right)\right)$ (4)
and since $\|Y\|_{1}\leq r$,
@@
$\leq\exp\left(t\left(-\frac{4\pi^{2}\sigma^{2}}{nq^{2}}\left(1+\Gamma_{ij}\right)+\frac{\mathbb{E}((\frac{2\pi(\delta_{i}+\delta_{j})}{q}\cdot Y)^{4})}{24}\right)\right)$
@@
By a similar argument (5),
$\exp\left(t\left(\Phi_{Y}\left(\frac{2\pi(\delta_{i}-\delta_{j})}{q}\right)-1\right)\right)\leq\exp\left(-t\left(\frac{4\pi^{2}\sigma^{2}}{nq^{2}}\right)\left(1-\sup_{k\neq\ell}|\Gamma_{k\ell}|\right)\left(1-\frac{\pi^{2}r^{2}}{3q^{2}}\right)\right).$ (6)
@@
$\operatorname{Var}_{\nu_{t}}(\psi)=\sum_{i=1}^{n}\operatorname{Var}_{\nu_{t}}\left(\cos\left(\frac{2\pi X_{t}(i)}{q}\right)\right)+\sum_{i=1}^{n}\sum_{j\neq i}\operatorname{Cov}_{\nu_{t}}\left(\cos\left(\frac{2\pi X_{t}(i)}{q}\right),\cos\left(\frac{2\pi X_{t}(j)}{q}\right)\right)$ (9)
@@
where the last expression vanishes to 00 as $n\to\infty$ by the assumptions (1). Hence, for sufficiently large $n$,
@@ 4.1 Introduction
Recall from Section 1 the dynamics of the rate-$1$ continuous-time random walk over the $1\times n$ contingency tables with any fixed coordinate sum. We will only keep track of the first $n-1$ coordinates so that the resulting process $\{X_{t}\}_{t\geq 0}$ on ${\mathbb{Z}}_{q}^{n-1}$ will be aperiodic and irreducible. We will use $({\mathbb{Z}}\cap[-q/2,q/2))^{n-1}$ as representatives of ${\mathbb{Z}}_{q}^{n-1}$. Let $\{Y_{m}\}_{m\in{\mathbb{N}}}$ be the embedded discrete-time chain of $\{X_{t}\}$. Since our processes are vertex-transitive, without loss of generality, we may assume that $X_{0}=0$ and $Y_{0}=0$. Denote by $P$ the transition matrix of $\{Y_{m}\}$. Then the rate matrix $Q$ for $\{X_{t}\}$ is $Q=P-I$ where $I$ is the identity matrix, and the transition kernel $H_{t}$ for $\{X_{t}\}$ is $H_{t}=\exp(tQ)$. We write
@@
where $a_{ij}\in{\mathbb{Z}}_{q}^{n-1}$ is defined as follows: if $11\leq i,j\leq n-$ 1, then $a_{ij}$ has $1$ on the $i$-th coordinate, $-1$ on the $j$-th coordinate, and zero elsewhere; if $i=n$, then $a_{ij}$ has $-1$ on the $j$-th coordinate and zero elsewhere; if $j=n$, then $a_{ij}$ has $1$ on the $i$-th coordinate and zero elsewhere. Then $\{Y_{m}\}$ is a random walk on the group ${\mathbb{Z}}_{q}^{n-1}$ with the symmetric set of generators $S$.
@@
4.2 Eigenvalues for $\{X_{t}\}$ and $\{Y_{m}\}$
By the same reasoning as in Section 3.2, the Fourier transforms $\{\hat{\rho}_{y}:y\in{\mathbb{Z}}_{q}^{n-1}\}=\{\Phi_{Y}(2\pi y/q):y\in({\mathbb{Z}}\cap[-q/2,q/2))^{n-1}\}$ are precisely the eigenvalues of the discrete-time process $\{Y_{m}\}$, where $\Phi_{Y}$ is again the characteristic function of $Y$. And the eigenvalues $\{\lambda_{j}:1\leq j\leq q^{n-1}\}$ of the continuous-time process $\{X_{t}\}$ are
$\{\lambda_{j}:1\leq j\leq q^{n-1}\}=\{\exp(t(\Phi_{Y}(2\pi y/q)-1)):y\in({\mathbb{Z}}\cap[-q/2,q/2))^{n-1}\}.$

### v2

@@ 2.2 Technical Lemmas
We break the sum into two parts and bound them, respectively, as follows. We can write
@@
If $\kappa_{1}n\leq m\leq\kappa_{2}n$ with $0<\kappa_{1}\leq 1\leq\kappa_{2}$ and $\geq 1, mm\in{\mathbb{N}}$, then
$\sum_{\ell=1}^{m}\sum_{k=1}^{n}\left(\frac{1}{mn}\right)^{\frac{\ell}{\ell+1}\frac{k}{k+1}}\leq\frac{\kappa_{2}^{2}(1+2e^{4/e}+2e^{8/e}+e^{32/e})}{\kappa_{1}}.$
@@
We will prove the first bound in the claim, and the second bound will follow as a straightforward corollary. For any $n\in\mathbb{N}$, we can write
@@
$\frac{1}{n^{2}}\sum_{\ell=1}^{\lfloor n^{1/4}\rfloor}\sum_{k=1}^{\lfloor n^{1/4}\rfloor}n^{\frac{2(k+\ell+1)}{(k+1)(\ell+1)}}\leq\frac{1}{n^{2}}\cdot n^{1/4}\cdot n^{1/4}\cdot n^{3/2}=1.$ (1)
@@
$\frac{1}{n^{2}}\sum_{\ell=\lfloor n^{1/4}\rfloor}^{n}\sum_{k=\lfloor n^{1/4}\rfloor}^{n}n^{\frac{2(k+\ell+1)}{(k+1)(\ell+1)}}\leq\frac{1}{n^{2}}\sum_{\ell=\lfloor n^{1/4}\rfloor}^{n}\sum_{k=\lfloor n^{1/4}\rfloor}^{n}n^{\frac{4(k+\ell)}{\ell k}}\leq\frac{1}{n^{2}}\cdot n\cdot n\cdot n^{8n^{-1/4}}\leq e^{32/e}$ (2)
as $x^{8x^{-1/4}}$ is maximized as $e^{32/e}$ at $x=e^{4}$ by elementary calculus. Now it suffices to bound the sum for $\ell\leq n^{1/4}$ and $k\geq n^{1/4}$, and the cases $\ell\geq n^{1/4}$ and $k\leq n^{1/4}$ would follow in an analogous way. We consider the scenario of $\ell=1$ separately. When $\ell=1$, we have
$\frac{1}{n^{2}}\sum_{k=\lfloor n^{1/4}\rfloor}^{n}n^{\frac{k+2}{k+1}}=\frac{1}{n}\sum_{k=\lfloor n^{1/4}\rfloor}^{n}n^{\frac{1}{k+1}}\leq\frac{1}{n}\cdot n\cdot n^{n^{-1/4}}\leq e^{4/e}.$ (3)
@@
$\frac{1}{n^{2}}\sum_{\ell=2}^{\lfloor n^{1/4}\rfloor}\sum_{k=\lfloor n^{1/4}\rfloor}^{n}n^{\frac{2(k+\ell+1)}{(k+1)(\ell+1)}}\leq\frac{1}{n^{2}}\cdot n^{1/4}\cdot n\cdot n^{\frac{2(n^{1/4}+3)}{3n^{1/4}}}=n^{-1/12}\cdot n^{2n^{-1/4}}\leq e^{8/e}.$ (4)
Combining (1), (2), (3), and (4) establishes the first bound in the claim.
Now for the second bound, we can write
$\sum_{\ell=1}^{m}\sum_{k=1}^{n}\left(\frac{1}{mn}\right)^{\frac{\ell}{\ell+1}\frac{k}{k+1}}\leq\sum_{\ell=1}^{\kappa_{2}n}\sum_{k=1}^{\kappa_{2}n}\left(\frac{\kappa_{2}^{2}n/m}{(\kappa_{2}n)^{2}}\right)^{\frac{\ell}{\ell+1}\frac{k}{k+1}}\leq\frac{\kappa_{2}^{2}n}{m}\sum_{\ell=1}^{\kappa_{2}n}\sum_{k=1}^{\kappa_{2}n}\left(\frac{1}{(\kappa_{2}n)^{2}}\right)^{\frac{\ell}{\ell+1}\frac{k}{k+1}}.$
Since $m\geq\kappa_{1}n$, $\dfrac{\kappa_{2}^{2}n}{m}\leq\dfrac{\kappa_{2}^{2}}{\kappa_{1}}$ and the result follows from the first bound. ∎
We note that in Lemmas 2, 3, the constants $e^{2/e},e^{4/e},$ etc. are chosen for convenience to give explicit uniform bounds.
@@
Let $F(x):=\sum_{j=-N}^{N}\exp(-c(j-x)^{2})$, where $N\in{\mathbb{N}}$ and $c>0$ are constants. Then there exists a constant $c_{0}>0$ independent of $N$, such that whenever $c>c_{0}$, $F(x)$ attains its global maximum at $x=0$.
@@
First, we observe that for any $x\in(-1/2,1/2]$, we have
@@
by comparing the summands in $F(x+k)$ and $F(x+k+1)$, $k\in{\mathbb{Z}}$. Thus, it suffices to maximize $F$ on the interval (1/2,1/2](-1/2,1/2].
@@
We will only consider $c>0$ that is not too small. If $x\not\in\bigcup_{j=-N}^{N}B_{j}(1/\sqrt{2c})$, then $F^{\prime\prime}(x)$ is necessarily positive, since all the summands above are positive. Combining this with the previous observation, we only need to maximize $F$ on the interval $1/\sqrt{2c},1/\sqrt{2c}]$.
@@
Note that for all $c$ not too small and any fixed $x\in[0,1/\sqrt{2c}]$, the function
@@
Note that for all $c$ not too small, the function $g(x):=(1+x+\frac{1}{2c})e^{-c(1+x)^{2}}$ is convex on the interval $1/\sqrt{2c},1/\sqrt{2c}]$. So for $x\in[0,1/\sqrt{2c}]$,
@@
Now, it remains to show that
@@ 3.1 Setup
We provide a general method for mixing time lower bounds for a family of random walks on the torus ${\mathbb{Z}}_{q}^{n}$. We parameterize the torus ${\mathbb{Z}}_{q}^{n}$ by $({\mathbb{Z}}\cap[-q/2,q/2))^{n}$, i.e. these are the representatives we will use for elements in ${\mathbb{Z}}_{q}^{n}$.
Let $\{Y_{s}\}_{s\in{\mathbb{N}}}$ be an arbitrary irreducible symmetric discrete-time Markov chain on ${\mathbb{Z}}^{n}_{q}$ with transition matrix $P$, such that ${\mathbb{P}}(Y_{s}=y_{s}|Y_{s-1}=y_{s-1})=\mu(y_{s}-y_{s-1})$. Suppose that $\|Y\|_{1}\leq r$ for some real number $r$ in the $L_{1}$ norm, and $\|Y\|_{\infty}<q/2$ in the sup norm. We set $Y_{0}=0$. Symmetry of $\{Y_{s}\}$ implies $\mu(y)=\mu(-y)$ for all $y\in{\mathbb{Z}}_{q}^{n}$. Let $Y\in{\mathbb{Z}}^{n}$ be a lattice-valued random vector with law $\mu$ mapped to $({\mathbb{Z}}\cap[-q/2,q/2))^{n}$. We will assume $Y$ has the same marginal variance $\dfrac{\sigma^{2}}{n}$ for each coordinate. Write $\operatorname{Var}(Y)=\dfrac{\sigma^{2}}{n}\Gamma$ where $\Gamma$ is the correlation matrix of $Y$. Let $\{X_{t}\}_{t\geq 0}$ be the rate-1 continuous-time Markov chain on ${\mathbb{Z}}_{q}^{n}$ with rate matrix Q=PIQ=P-I and transition kernel $H_{t}=\exp(tQ)$. Set $X_{0}=0$.
@@ 3.2 Representations and Eigenvalues
By Theorem 3 of [10] (a more detailed proof could be found in Theorem 6 of [6]), the Fourier transforms $\{\hat{\mu}(\rho_{y}):y\in{\mathbb{Z}}_{q}^{n}\}$ are precisely the eigenvalues of the discrete-time process $\{Y_{s}\}$, and the characters are the eigenfunctions of $P$. Thus the eigenvalues $\{\lambda_{j}:1\leq j\leq q^{n}\}$ of the continuous-time process $\{X_{t}\}$ are
$\{\lambda_{j}:1\leq j\leq q^{n}\}=\{\exp(\Phi_{Y}(2\pi y/q)-1):y\in({\mathbb{Z}}\cap[-q/2,q/2))^{n}\}.$
@@ 3.3.1 Forming a Distinguishing Statistic
Let
$\psi_{i}(x):=\cos\left(\dfrac{2\pi x_{i}}{q}\right)$
for $({\mathbb{Z}}\cap[-q/2,q/2))^{n}$. We form a distinguishing statistic
@@
Let $\delta_{i}$ be the vector whose $i$-th coordinate is $1$ and zero elsewhere for notation. We know that each $\psi_{i}$ is an eigenfunction of $H_{t}$ with a corresponding eigenvalue of
@@
By the fourth-order Taylor expansion of the characteristic function $\Phi_{Y}$, we obtain that for any $i$,
$\exp\left(-\frac{2\pi^{2}t\sigma^{2}}{nq^{2}}\right)\leq\exp\left(t\left(\Phi_{Y}\left(\frac{2\pi\delta_{i}}{q}\right)-1\right)\right).$ (8)
The fourth-order Taylor expansion of $\Phi_{Y}\left(\dfrac{2\pi(\delta_{i}+\delta_{j})}{q}\right)$ for $i\neq j$ gives
@@
$\leq\exp\left(t\left(-\dfrac{2\pi^{2}\sigma^{2}}{nq^{2}}\left(\delta_{i}+\delta_{j}\right)^{T}\Gamma\left(\delta_{i}+\delta_{j}\right)+\frac{\mathbb{E}((\frac{2\pi(\delta_{i}+\delta_{j})}{q}\cdot Y)^{4})}{24}\right)\right)$
$=\exp\left(t\left(-\frac{4\pi^{2}\sigma^{2}}{nq^{2}}\left(1+\Gamma_{ij}\right)+\frac{\mathbb{E}((\frac{2\pi(\delta_{i}+\delta_{j})}{q}\cdot Y)^{4})}{24}\right)\right)$
Recall the assumption $\|Y\|_{1}\leq r$, so by Cauchy-Schwarz, the above is upper bounded by
@@
By an argument similar to (9),
$\exp\left(t\left(\Phi_{Y}\left(\frac{2\pi(\delta_{i}-\delta_{j})}{q}\right)-1\right)\right)$ (10)
@@
$\operatorname{Var}_{\nu_{t}}(\psi)$ $=\sum_{i=1}^{n}\operatorname{Var}_{\nu_{t}}\left(\cos\left(\frac{2\pi X_{t}(i)}{q}\right)\right)$ (13)
$+\sum_{i=1}^{n}\sum_{j\neq i}\operatorname{Cov}_{\nu_{t}}\left(\cos\left(\frac{2\pi X_{t}(i)}{q}\right),\cos\left(\frac{2\pi X_{t}(j)}{q}\right)\right)$
@@
where the last expression vanishes to 0 as $n\to\infty$ by the assumptions (5). Hence, for sufficiently large $n$,
@@ 4.1 Introduction
Recall from Section 1 the dynamics of the rate-$1$ continuous-time random walk over the $1\times n$ contingency tables with any fixed coordinate sum. We will only keep track of the first n1n-1 coordinates so that the resulting process $\{X_{t}\}_{t\geq 0}$ on ${\mathbb{Z}}_{q}^{n-1}$ will be aperiodic and irreducible. We will use $({\mathbb{Z}}\cap[-q/2,q/2))^{n-1}$ as representatives of ${\mathbb{Z}}_{q}^{n-1}$. Let $\{Y_{s}\}_{s\in{\mathbb{N}}}$ be the embedded discrete-time chain of $\{X_{t}\}$. Since our processes are vertex-transitive, without loss of generality, we may assume that $X_{0}=0$ and $Y_{0}=0$. Denote by $P$ the transition matrix of $\{Y_{s}\}$. Then the rate matrix $Q$ for $\{X_{t}\}$ is Q=PIQ=P-I where $I$ is the identity matrix, and the transition kernel $H_{t}$ for $\{X_{t}\}$ is $H_{t}=\exp(tQ)$. We write
@@
where $a_{ij}\in{\mathbb{Z}}_{q}^{n-1}$ is defined as follows: if $1\leq i,j\leq n-1$, then $a_{ij}$ has $1$ on the $i$-th coordinate, 1-1 on the $j$-th coordinate, and zero elsewhere; if $i=n$, then $a_{ij}$ has 1-1 on the $j$-th coordinate and zero elsewhere; if $j=n$, then $a_{ij}$ has $1$ on the $i$-th coordinate and zero elsewhere. Then $\{Y_{s}\}$ is a random walk on the group ${\mathbb{Z}}_{q}^{n-1}$ with the symmetric set of generators $S$.
@@
4.2 Eigenvalues for $\{X_{t}\}$ and $\{Y_{s}\}$
By the same reasoning as in Section 3.2, the Fourier transforms $\{\hat{\rho}_{y}:y\in{\mathbb{Z}}_{q}^{n-1}\}=\{\Phi_{Y}(2\pi y/q):y\in({\mathbb{Z}}\cap[-q/2,q/2))^{n-1}\}$ are precisely the eigenvalues of the discrete-time process $\{Y_{s}\}$, where $\Phi_{Y}$ is again the characteristic function of $Y$. And the eigenvalues $\{\lambda_{j}:1\leq j\leq q^{n-1}\}$ of the continuous-time process $\{X_{t}\}$ are
$\{\lambda_{j}:1\leq j\leq q^{n-1}\}=\{\exp(\Phi_{Y}(2\pi y/q)-1):y\in({\mathbb{Z}}\cap[-q/2,q/2))^{n-1}\}.$